### v1

# Noise-Conditioned Mixture-of-Experts Framework for Robust Speaker Verification

Abstract
Robust speaker verification under noisy conditions remains an open challenge. Conventional deep learning methods learn a robust unified speaker representation space against diverse background noise and achieve significant improvement. In contrast, this paper presents a noise-conditioned mixture-of-experts framework that decomposes the feature space into specialized noise-aware subspaces for speaker verification. Specifically, we propose a noise-conditioned expert routing mechanism, a universal model based expert specialization strategy, and an SNR-decaying curriculum learning protocol, collectively improving model robustness and generalization under diverse noise conditions. The proposed method can automatically route inputs to expert networks based on noise information derived from the inputs, where each expert targets distinct noise characteristics while preserving speaker identity information. Comprehensive experiments demonstrate consistent superiority over baselines, confirming that explicit noise-dependent feature modeling significantly enhances robustness without sacrificing verification accuracy.
Index Terms— Speaker verification, mixture-of-experts, noise robustness.
1 Introduction
Speaker verification (SV), which aims to verify the identity of a given utterance[1], has been widely adopted in smart devices and other security-critical applications. While deep learning has substantially advanced SV systems [2, 3, 4, 5, 6], their real-world deployment still faces significant challenges, as the shift to unconstrained environments brings challenging acoustic interference [7]. Common noise sources, including ambient music, non-stationary noise, and crowd babble, create diverse spectral distortions that substantially degrade verification performance. This robustness challenge continues to hinder reliable real-world implementation.
To alleviate the above-mentioned problem, a dominant and effective approach is to employ speech enhancement (SE) networks for noise suppression, with the enhanced feature then utilized for SV [8, 9]. Compared to cascading pre-trained SE and SV networks, [10] showed that constructing an SE model specialized for SV tasks yields better performance. Thereafter, some studies have attempted to boost system robustness through targeted improvements to SE modules. In [11], both masking- and mapping-based SE networks are integrated into the SV system to remove noise from different aspects. In [12], a novel extened U-Net is adopted as the SE backbone and showed the superiority of the jointly optimized cascade system through end-to-end learning. With the advent of diffusion models, recent works [13][14] have incorporated them into the SE front-end, achieving notable performance gains. Despite these advances, the customized architectures of such cascaded systems grow prohibitively complex by incorporating both enhancement and verification processes, leading to computational inefficiency that hinders practical deployment. Moreover, error propagation remains a fundamental concern as spectral distortions introduced during enhancement may adversely impact downstream verification accuracy.
An alternative technical paradigm employs advanced learning strategies to derive noise-invariant speaker representations. As demonstrated in [15], feeding clean and noisy speech into the network while minimizing the distances at the embedding level can improve robustness. Further improvements are achieved by [14] through supervised contrastive learning to narrow distribution gaps between clean and noisy samples, and by [16] via disentanglement methods to isolate noise components from speech representations. [17] extends this framework by incorporating adversarial training to treat different noise types as distinct domains, making their representations domain indistinguishable. Additionally, [18] introduces stable learning to eliminate spurious correlations in training data, thereby improving noise generalization. However, such representation learning methods typically involve multiple loss functions, increasing training complexity and requiring careful tuning of loss weights for optimal performance.
While both kinds of approaches have shown promising results, their dependence on unified feature-space modeling may present certain constraints. In cases where input distributions exhibit notable variations, maintaining effective discrimination within a single feature space could prove challenging. Additionally, the strategy of aligning clean and noisy embeddings might inadvertently affect voiceprint fidelity, as the optimization process could nudge pristine features toward noise-adapted positions. This potential trade-off between noise robustness and speaker discriminability suggests opportunities for further refinement, particularly for handling extreme acoustic variability.
Based on above analysis, we explore a noise-conditioned mixture-of-experts (NCMoE) framework, which has gained significant traction recently in natural language processing [19][20][21]. Rather than relying on a unified feature space, our method investigates the potential of decomposing the feature space into noise-specific subspaces, with the aim of improving speaker feature extraction in challenging acoustic environments. Firstly, a lightweight convolutional network is employed to estimate the noise characteristics of spectral features, which then guides the selection of specialized expert branches. To facilitate expert training, then we propose a universal model based learning strategy which begins with a generalist expert model before progressively specializing in noise-adapted subspaces, complemented by an SNR-progressive curriculum that appears to enhance training stability. Experiments on the VoxCeleb1 dataset with simulated noise conditions suggest that this approach performs better than existing methods. We hope that this exploration of condition-specific subspace modeling might contribute to ongoing discussions about robust speaker representation learning.
2 Noise-Conditioned Mixture-of-Expert Framework
2.1 Overview of the Framework
As illustrated in Fig.1, our framework preserves the original ResNet architecture while augmenting its second residual stage with parallel expert branches to balance model capacity and efficiency. Each expert replicates the second stage’s complete residual structure (i.e. the sequence of residual blocks with identical channel dimensions and skip connections), maintaining architectural consistency with the backbone. In addition, a compact noise classification network (Fig.1, bottom-right) dynamically selects a single expert branch per sample during forward propagation, keeping others inactive. This design ensures computational efficiency through sparse expert activation while maintaining manageable resource requirements, allowing specialized processing tailored to distinct noise conditions and retains the ResNet backbone’s native computational patterns.
2.2 Noise-Conditioned Expert Routing
The Noise-Conditioned Expert Routing (NCER) method dynamically selects processing paths based on input features. Given an input feature , the noise classifier predicts the noise category distribution , where each routing value for routing expert is computed via temperature-scaled softmax:
The classifier consists of sequential strided convolutions followed by temporal pooling and softmax, with denoting the logit for noise class and temperature factor controlling output sharpness. The combination of different experts is calculated as:
where are expert networks. The conditional execution strategy in Eq.2 ensures proper gradient flow through all experts during backpropagation (training phase), while eliminating unnecessary computational overhead from inactive branches during inference (testing phase). Note that all experts share identical architectures, and the noise classifier uses only three convolutional layers. This simplicity ensures performance gains stem solely from the routing strategy, not network architecture design.
2.3 Universal Model Based Expert Specialization
Inspired by the GMM-UBM [22], we employ a Universal Model based Expert Specialization (UMES) strategy where expert networks first learn shared feature representations before specializing. The entire training process is evenly divided into two distinct phases based on the total number of epochs, ensuring a clear transition from shared learning to specialized adaptation. During Phase I, all expert networks process identical input data using the same initialized parameters , the optimization objective is formulated as:
where optimizes the noise classifier while optimizes the speaker classifier based on the expert combination function . Through uniform output aggregation, all experts consequently receive identical parameter updates during backpropagation, which is effectively equivalent to first training a universal model and then replicating it into multiple experts. The optimization objective in Phase II is formulated as:
where each expert develops distinct parameters through -weighted gradient updates , and thus noise-conditional specialization is introduced. In addition, by retaining the Phase I loss, the Phase II objective ensures robust speaker recognition across all conditions while facilitating expert-specific specialization.
2.4 SNR-Decaying Curriculum Learning
The SNR-Decaying Curriculum Learning (SDCL) enhances model learning efficiency through progressive SNR reduction in training data augmentation. The method implements an easy-to-hard curriculum where the augmentation SNR follows a truncated Gaussian distribution:
where denotes a truncated Gaussian distribution that bounds the SNR within meaningful limits to avoid ineffective SNR levels, represents the epoch-dependent mean that controls curriculum progression, and determines the sampling variability around the mean to maintain training diversity. The mean SNR decays exponentially across training epochs according to:
where is the decay rate coefficient, indicates the current training epoch, and represents the total number of training epochs that normalize the progression.
Consequently, the SDCL prevents early exposure to extreme noise conditions that could hinder convergence, while simultaneously enabling gradual adaptation to diverse SNR levels. Moreover, it effectively facilitates expert specialization across different SNR ranges, and through its controlled noise introduction it consistently maintains training stability throughout the learning process. Finally, the truncated Gaussian sampling provides beneficial stochastic variability while strictly respecting the curriculum progression framework.
3 Experiments
3.1 Data
The experiments are conducted on the standard development and test sets of the Voxceleb1 dataset [23]. The development set contains 1211 speakers for training, while the test set consists of 37720 trials from 40 speakers for evaluation. To thoroughly assess system robustness, we construct noisy test conditions by augmenting the original clean utterances with MUSAN [24] and Nonspeech100 [25] at signal-to-noise ratios (SNRs) ranging from 0 to 20 dB in 5 dB increments. The MUSAN dataset includes “babble”, “music” and “noise”1 as three distinct audio types, which are partitioned into non-overlapping training and testing subsets to prevent data leakage, following the partition protocol in [13]. During model training, Each utterance undergoes online data augmentation through either additive noise mixing (with randomly selected samples from the MUSAN training set) or convolutional reverberation (using simulated room impulse responses). The remaining noise samples are reserved for synthesizing comprehensive evaluation sets that cover both in-domain and out-of-domain noise conditions.
3.2 Implementation Details
The acoustic front-end extracts 80-dimensional log-mel filterbank features. The experiments are based on the ResNet34 backbone with 32 initial channels, and the embedding dimension is 256. The model is trained for 150 epochs using SGD optimizer with the AAM-Softmax loss for speaker classification, employing mixed precision training to accelerate computation while maintaining stability. The proposed framework extends this baseline with four parallel residual branches in residual stage two, each following the identical structure as mentioned in Section II-2.1, while the noise classifier maintains a 32-channel initial width and adheres to the protocol of doubling channels during time-frequency downsampling. We treat reverberation as a structured noise and accordingly define four distinct augmentation categories as noise-type labels: “noise”, “babble”, “music” and “reverberation”. The temperature factor is empirically set at 0.1 , with the truncated Gaussian distribution bounded between 20 dB and 0 dB, and its mean empirically set at 0.2. The coefficient in Eq.6 is derived as 7.6 by logarithmic transformation to ensure that decays smoothly from 20 to 0.01 (approximately 0 for numerical stability) during training. The implementation is based on the Wespeaker [26] toolkit using 2 NVIDIA RTX 5070 Ti GPUs.
3.3 Results
| Training Set | VoxCeleb1 | |||||||||||
| Noise Type | SNR | Baseline | VoiceID[10] | FSEF[11] | NDML[16] | WSVIL[15] | ExU-Net[12] | SEU-Net[18] | Diff-SV[13] | NDAL[17] | NISRL[14] | NCMoE |
| Original Set | 1.98 | 6.79 | 4.26 | 2.90 | 3.12 | 2.76 | 2.52 | 2.35 | 2.63 | 2.40 | 1.91 | |
| Babble | 0 | 9.30 | 38.0 | 27.6 | 11.0 | 11.8 | 9.57 | 8.54 | 8.74 | 6.43 | 7.81 | 8.10 |
| 5 | 4.56 | 27.1 | 15.3 | 6.13 | 5.97 | 5.52 | 5.16 | 4.51 | 4.44 | 4.25 | 4.24 | |
| 10 | 2.99 | 16.7 | 9.04 | 4.28 | 4.44 | 4.06 | 3.67 | 3.33 | 3.59 | 3.28 | 2.88 | |
| 15 | 2.45 | 11.3 | 6.47 | 3.52 | 3.73 | 3.28 | 3.10 | 2.82 | 3.08 | 2.78 | 2.51 | |
| 20 | 2.18 | 8.99 | 5.41 | 3.21 | 3.36 | 2.99 | 2.79 | 2.61 | 2.87 | 2.60 | 2.06 | |
| Music | 0 | 5.82 | 16.2 | 8.47 | 10.8 | 7.79 | 7.35 | 6.25 | 6.04 | 5.87 | 5.19 | 4.62 |
| 5 | 3.57 | 11.4 | 6.31 | 6.52 | 5.23 | 4.90 | 4.36 | 3.96 | 4.19 | 3.58 | 3.04 | |
| 10 | 2.73 | 9.13 | 5.14 | 4.66 | 4.11 | 3.69 | 3.55 | 3.10 | 3.53 | 3.11 | 2.50 | |
| 15 | 2.28 | 8.10 | 4.71 | 3.67 | 3.63 | 3.14 | 3.10 | 2.75 | 3.23 | 2.75 | 2.11 | |
| 20 | 2.13 | 7.48 | 4.56 | 3.21 | 3.30 | 2.93 | 2.79 | 2.60 | 3.09 | 2.57 | 2.07 | |
| Noise | 0 | 7.3 | 16.6 | 7.88 | 10.2 | 7.34 | 6.80 | 6.41 | 6.01 | 6.14 | 4.94 | 5.20 |
| 5 | 4.45 | 12.3 | 6.42 | 6.96 | 5.65 | 5.23 | 4.42 | 4.52 | 4.00 | 3.69 | 3.65 | |
| 10 | 3.14 | 9.86 | 5.50 | 5.02 | 4.35 | 4.07 | 3.74 | 3.49 | 3.23 | 3.43 | 2.72 | |
| 15 | 2.57 | 8.69 | 4.87 | 3.91 | 3.85 | 3.39 | 3.20 | 2.93 | 2.97 | 2.94 | 2.39 | |
| 20 | 2.25 | 7.83 | 4.66 | 3.40 | 3.44 | 3.10 | 2.92 | 2.64 | 2.80 | 2.68 | 2.16 | |
| Average | 3.73 | 13.5 | 7.91 | 5.59 | 5.07 | 4.55 | 4.16 | 3.90 | 3.88 | 3.62 | 3.26 | |
| SNR | Baseline | SEU-Net | Diff-sv | NDAL | NISRL | NCMoE |
| 0 | 10.17 | 5.99 | 8.23 | 7.57 | 6.41 | 6.27 |
| 5 | 5.53 | 4.58 | 5.06 | 5.49 | 4.57 | 4.02 |
| 10 | 3.79 | 3.74 | 3.85 | 4.03 | 3.55 | 3.05 |
| 15 | 2.74 | 3.15 | 3.19 | 3.36 | 2.99 | 2.43 |
| 20 | 2.36 | 2.87 | 2.89 | 2.99 | 2.75 | 2.19 |
| Average | 4.92 | 4.07 | 4.65 | 4.97 | 4.05 | 3.59 |
Table 1 presents a comparative evaluation between the proposed method and existing approaches. Longitudinally, performance degrades under all simulated noisy conditions compared to the original test set, with particularly severe degradation at SNRs below 10 dB. Among the three noise types (music, babble, and non-stationary noise), the most pronounced degradation occurs under babble noise, while music noise exhibits relatively milder effects. We attribute this to the closer spectral similarity between babble and voiceprint information, which amplifies interference with speaker characteristics, whereas music’s distinct spectral patterns facilitate easier separation of speaker features.
Horizontally, the proposed method demonstrates significant improvements over baseline systems and competing approaches, achieving state-of-the-art results. The baseline itself outperforms most comparative methods, which we hypothesize stems from the advanced learning configurations [26]. Compared to the two kinds of methodologies outlined in Section I, performance varies across methods, with NPSRL (which hybridizes both paradigms) delivering the best results among prior work. Nevertheless, our method surpasses even NPSRL, validating the superiority of independent modeling in conditioned noise spaces. Notably, the average performance across noise conditions does not strictly correlate with performance on the clean dataset among different comparison systems, suggesting that superior performance under specific noise types may not generalize to pristine conditions. This observation underscores the need for comprehensive multi-condition testing to validate algorithmic robustness. Further analysis in Table 2 evaluates cross-domain generalization under unseen noise conditions. The proposed method exhibits consistent performance gains, confirming its enhanced robustness in unknown scenarios.
3.4 Ablation Study and Analysis
| Noise Type | NCMoE | w/o UMES | w/o NCER | w/o SDCL |
| Babble | 3.96 | 9.50 | 4.02 | 4.05 |
| Music | 2.87 | 7.14 | 3.00 | 2.98 |
| Noise | 3.23 | 8.37 | 3.43 | 3.30 |
| Nonspeech | 3.59 | 8.99 | 3.96 | 3.81 |
| Average | 3.41 | 6.80 | 3.60 | 3.54 |
Our ablation study in Table 3 demonstrates varying degrees of performance degradation when removing three key components: the UMES, NCER, and SDCL. The most significant performance drop occurs with the removal of the UMES strategy, indicating that excessive specialization among experts compromises discriminative speaker feature extraction due to their inherent shared patterns. Although eliminating the noise classification loss (w/o NCER) degrades performance, the model still maintains superiority over the baseline, revealing both its intrinsic sample routing capability and the additional benefit of explicit noise-type supervision for condition-specific modeling. Furthermore, the ablation of SDCL confirms that progressively reducing SNR during training effectively enhances model robustness, establishing the value of phased difficulty escalation in the learning process.
4 Conclusion
In this work, we propose a noise-conditioned mixture-of-experts framework for robust speaker verification. Unlike existing approaches that rely on a robust unified feature modeling space, our method decomposes the feature forwarding space into multiple noise-specific subspaces, enabling superior handling of diverse noise conditions. Extensive experiments validate the effectiveness of our proposed framework, demonstrating consistent performance improvements over competing methods. Future work will focus on investigating more advanced mixture architectures, refined noise classifiers, and broader noise categorizations to further enhance the model’s capabilities.